OpenStudy (superhelp101):

When one atom of carbon (C) combines with four atoms of hydrogen (H), four nonpolar covalent bonds form, resulting in a molecule of methane (CH4). Which of the following describes the arrangement of valence electrons in a bond between C and H?

OpenStudy (superhelp101):

The valence electrons spend more time around the atom of C. The valence electrons spend more time around the atom of H. The valence electrons are shared equally between C and H. The valence electrons are given up by the H and gained by the C.

OpenStudy (abhisar):

What do u think it should be ?

OpenStudy (superhelp101):

the 3rd option???

OpenStudy (abhisar):

Why ?

OpenStudy (superhelp101):

opps i meant the 2 one

OpenStudy (abhisar):

Random guessing ??

OpenStudy (abhisar):

What are the electronegativities of hydrogen and carbon ?

OpenStudy (superhelp101):

carbon is 2.5 and hydrogen is 2.1

OpenStudy (superhelp101):

is it right?

OpenStudy (abhisar):

Who will attract the shared electrons more ?

OpenStudy (superhelp101):

hydrogen

OpenStudy (abhisar):

The element which has higher electronegativity attracts the shared electron pair more.

OpenStudy (abhisar):

Carbon has > EN than hydrogen so it will attract the shared electron pair more.

OpenStudy (superhelp101):

oh

OpenStudy (abhisar):

So the correct option will be ??

OpenStudy (superhelp101):

second

OpenStudy (abhisar):

How ??? Carbon will attract the shared pair more, so it will spend more time around carbon.

OpenStudy (abhisar):

getting my points ?

OpenStudy (superhelp101):

yeah it the first one

OpenStudy (abhisar):

Yep, but then why it's non polar ?

OpenStudy (superhelp101):

because it contains one element

OpenStudy (abhisar):



OpenStudy (abhisar):

It's due to its tetrahedral geometry. Net dipole moment becomes zero

OpenStudy (superhelp101):

okay, alright that does make sentence. Thank you!!

OpenStudy (abhisar):

Sentence ?

OpenStudy (abhisar):

@Abmon98 need ur opinion !

OpenStudy (superhelp101):

Sense I meant lol

OpenStudy (abmon98):

If the atoms are equally electronegative, both have the same tendency to attract the bonding pair of electrons, and so it will be found on average half way between the two atoms.

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar but electronegativity of H and C is diff \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abmon98):

Let's apply the first step to CH4. The electronegativity of Carbon is 2.55 The electronegativity of Hydrogen is 2.20 Electronegativity difference between C and H = 2.55-2.20= 0.35 (The electronegativity difference ranging from 0.0 - 0.5 is considered to be negligible. Thus, non-polar.)

OpenStudy (abhisar):

So non-polarity is nothing to do with dipole moment ?

OpenStudy (abmon98):

yes :)

OpenStudy (abhisar):

u mean they have no relation ?

OpenStudy (abhisar):

yes is confusing me !

OpenStudy (abmon98):

Dipole Moment is related to polar molecule

OpenStudy (abhisar):

yeah i got uh now, actually question seeks bond polarity...ryt ??

OpenStudy (abmon98):

yes thats right

OpenStudy (abhisar):

@Abmon98 what about CCl bond ?

OpenStudy (abhisar):

It will have polar bond...ryt ?

OpenStudy (abhisar):

Electronegativity diff in CCl bond is 3-2.5 = 0.5, so it should be something like barely polar ryt ? @Abmon98

OpenStudy (abmon98):

Its Polar, polarity results in unequal sharing of valence electrons due to electronegativity therefore electrons will be more often located around the chlorine atom. this makes it a bit negative and electron cloud would spend more time and yes as you said due to small difference molecule is considered to be barely polar

OpenStudy (abhisar):

yep, because upto 0.5 it's considered non-polar.

OpenStudy (abmon98):

If the electronegativity difference is less than 0.5, then the bond is nonpolar covalent.

OpenStudy (abhisar):

ryt !

OpenStudy (abhisar):

not upto..it should be less than 0.5

OpenStudy (abhisar):

Thanx for the help @Abmon98 ! \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (abmon98):

Your most welcome :)

OpenStudy (superhelp101):

@abmon98 thanks for your help but I am really confused to which option it is according to your explanation.

OpenStudy (abhisar):

Check this @superhelp101 http://openstudy.com/updates/53ba5410e4b09f140ca6f6cd

OpenStudy (superhelp101):

yes. so it should be the second option right?

OpenStudy (superhelp101):

@Abhisar

OpenStudy (abhisar):

\(\bf Update:\) Electronegativity of carbon and hydrogen are 2.55 and 2.20 respectively which gives an electronegativity difference of 0.35. Any bond having a EN difference less than 0.4 is non-polar and hence electrons between hydrogen and carbon are shared equally. So the correct answer to the question on top should be option C Reference: http://www.chm.davidson.edu/ronutt/che115/electroneg.htm http://preparatorychemistry.com/Bishop_molecular_polarity.htm